If P(A or B) = 0.5, P(A) = 0.3 and P(B)= 0.2, determine
P(A and B).

Answer by ccs2011(207)

P(A or B) = P(A) + P(B) - P(A and B)
0.5 = 0.3 + 0.2 - P(A and B)
0.5 = 0.5 - P(A and B)
0 = -P(A and B)
0 = P(A and B)